What is the accumulated value of periodic deposits of $20 at the beginning of every quarter for 18 years if the interest rate is 3.45% compounded quarterly?
Answer #1

Calculation of accumulated value of quarterly deposit of $20 for 18 years in the following imageFuture value - PMT (1+r) -1 - - olkel &= Rate of interest = 3.45% = 3.45/100 = 0.0345 K- Payment forequency - Quarterly - 4 P

Accumulated value = $ 2001.64